40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Feb 22 07:53:06 2022, Yehonathan, Update, WOPO, Waking up WOPO 
    Reply  Fri Feb 25 17:09:53 2022, shruti, Update, WOPO, Waking up WOPO - green beam 
       Reply  Wed Mar 2 14:38:55 2022, Yehonathan, Update, WOPO, Waking up WOPO - some more fiddling and a plan 
          Reply  Fri Mar 4 11:49:45 2022, shruti, Update, WOPO, Waking up WOPO - attempts at readout LO_shot_noise.pdf
             Reply  Thu Mar 10 17:12:01 2022, awade, Update, WOPO, Waking up WOPO - attempts at readout 
                Reply  Wed Mar 16 12:22:44 2022, Yehonathan, Update, WOPO, Waking up WOPO - attempts at readout Moku_Noise.pdfPD_Connected_no_light.pdfDiff_Channel.pdf
                   Reply  Tue Mar 22 09:19:38 2022, shruti, Update, WOPO, Waking up WOPO - back to green path before_opt.jpgaround_waveguide.png
Message ID: 2725     Entry time: Fri Feb 25 17:09:53 2022     In reply to: 2723     Reply to this: 2727
Author: shruti 
Type: Update 
Category: WOPO 
Subject: Waking up WOPO - green beam 

[shruti, yehonathan]

SHG and 532 nm beam alignment

Yehonathan brought over 532nm/1064nm laser goggles from the 40m.

  • We turned on the 1064 nm Mephisto, and set the pump current initially to 1.5 A (which gave us ~30 mW of output)
  • We then turned on the doubling crystal. It took a while to reach its setpoint temperature of 110 C. Initially (without adjusting the SHG cavity parameters) we saw less than a microW of power after removing the beam block and opening the shutter.
  • Playing around with the SHG cavity settings, we realized that
    • Setting the switch to "Auto" is how to nominally operate the 532 nm beam
    • "Scan" is useful for scanning the cavity, the amplitude of which can be controlled by the "Scan amplitude" knob. "Standby" seems to effectively turn off second harmonic generation
    • "Offset" knob tends to change the amount of green power generated and adjusting the "Gain" simultaneously helps stabilize this power (with some difficulty)
  • On increasing the laser driver current to 2 A and adjusting the temperature setpoint to 109.7 C, we were able to see 100 mW of green power!
  • Next, we played with the alignment into the fiber, seeing that initially there was barely any power at the other end of that patch cable
    • We adjusted the waveplate near the laser head to give us 5.3 mW of power right before coupling into the fiber
    • Adjusting a single mirror (the nearer one) did not result in much gain, so we simultaneously adjusted the two steering mirrors and achieved about a 50% coupling. (Our readings suggested it could be the max)
    • At  the end of the alignment: 2.74 mW at the output and 5.46 at the fiber input
  • Reconnecting the fiber to the waveguide, without adjusting the temperature of the advr waveguide, we saw that the fiber at the output of this crystal seemed to glow.
    • The power measured at the end of the output fiber (fiber after the waveguide) was 0.6 mW. Not entirely sure what the contribution of loss was in the decrease from 2.74 mW through the waveguide.
  • The laser is still ON although the shutters to the green and IR paths are closed. Safety glasses required before opening shutters.

 

Questions about the setup

  1. The spec sheet on the wiki mentioned PM980 and PM480 input and output fibers, respectively for the waveguide operating as an SHG, what were being used instead were P3-1064PM-FC2 and P3-488PM-FC2. Is this a significant source of loss that can be easily remedied?
  2. Yehonathan mentioned that stimulated Brilluoin scattering occurs in all fibers above a threshold. What is the threshold for the the ones used in the setup? We probably want to operate below this threshold.

 

Our next step would be to measure the LO shot noise.

 

 

 

ELOG V3.1.3-